OpenStudy (anonymous):

Describe the differences between the graph of y = -3(x + 7)2 - 10 and the standard position graph of y = x2.

OpenStudy (anonymous):

y= x^2 looks like this

OpenStudy (anonymous):

The +7 does something, the -3 does something and the -10 does something to this graph

OpenStudy (anonymous):

so whats the difference?

OpenStudy (anonymous):

If you start with a graph of f(x), f(x+a) moves it a units to the left. f(x-a) moves it a units to the right. what does the 7 do to the graph? How does it shift the original graph?

OpenStudy (anonymous):

im too confused

OpenStudy (anonymous):

f(x) and y are the same thing If I start with y=x^2, the graph looks like the one I drew. What does y=(x+7)^2 look like? Its just the graph I drew shifted left 7 units. So the 7 inside the parenthesis in you problem will shift the parabola I drew 7 units to the left.

OpenStudy (anonymous):